6y^2-41y-7
( 6y-1)(y-7)



Answer by checkley71(8403)

6y~2-41y-7=0 or (6y+1)(y-7)=0 or 6y+1=0 or 6y=-1 or y=-1/6 and y-7=0 or y=7
6(-1/6)~2-41(-1/6)-7=0 or 6*1/36+41/6-7=0 or 6/36+41/6 -7=0 or 1/6+41/6-7=0 or
42/6-7=0 or 7-7=0